Before, solving the problem, I just want to give you the heads up. It's Application of Rational Equations. I can't find the specific title to relate with my story problems. My chapter in the book is Chapter Six Solving Rational Equations and Application of rational Equations. Hope it helps! (D=RxT)

*Elenore Morales traveled 1600 miles by commerical jet from Kansas City to Spokane. She then traveled as additional 500 miles on a private propeller plane from Spokane to Billings. If the speed of the jet was 4 times the speed of the propeller plane and the total time in the air was 6 hours, determine the speed of each plane.

Answer by stanbon(75887)

Elenore Morales traveled 1600 miles by commerical jet from Kansas City to Spokane.
She then traveled as additional 500 miles on a private propeller plane from Spokane to Billings.
If the speed of the jet was 4 times the speed of the propeller plane and the total time in the air was 6 hours, determine the speed of each plane.
------------------------------
KC to Spok DATA:
distance = 1600 miles ; rate = 4x mph ; time = 1600/4x = 400/x hrs
------------------------------
Spok to Bill DATA:
distance = 500 miles ; rate = x mph ; time = 500/x hrs
------------------------------
Equation:
time + time = 6 hrs
400/x + 500/x = 6
400 + 500 = 6x
6x = 900
x = 150 mph (speed of the propeller plane)
4x = 600 mph (speed of the jet plane)
========================================
Cheers,
Stan H.